1.   A college freshman must take a science course, a humanities course, and a mathematics course. If she may select any of 4 science courses, any of 3 humanities courses, and any of 2 mathematics courses, how many ways can she arrange her program?

2.   A clothing store wants to stock sweatshirts that come in four sizes (small, medium, large, x-large) and in three colors (yellow, blue, and white). How many different types of sweatshirts will the store have to stock?

3.   For her birthday, Karla received a new wardrobe consisting of 5 shirts, 3 pairs of pants, 2 skirts, and 3 pairs of shoes for her birthday. How many new outfits can she make?




1
Expert's answer
2022-05-04T12:44:13-0400

1.


"\dbinom{4}{1}\dbinom{3}{1}\dbinom{2}{1}=4(3)(2)=24"

2.


"\dbinom{4}{1}\dbinom{3}{1}=4(3)=12"

3.


"\dbinom{5}{1}\dbinom{3}{1}\dbinom{3}{1}+\dbinom{5}{1}\dbinom{2}{1}\dbinom{3}{1}"

"=5(3)(3)+5(2)(3)=75"